Q . FOR THE REACTION :-
   
   CaCO3 (s)------------ ------------> CaO (s) + CO2 (g),
                 Kp = 1.16 atm AT 8000C .
 
IF 20.0 g OF CaCO3  WERE KEPT IN A 10 L CONTAINER AND HEATED UP TO 8000 C , WHAT PERCENTAGE OF CaCO3 WOULD REMAIN UNREACTED AT EQUILIBRIUM.
                        

Grade:Select Grade

4 Answers

Sunil Kumar FP
9 years ago
we have Kp=(CO2)
mole of CaCO3 =weight of calcium carbonate/molcular weight
=20/100
=.2mole
In terms of molarity =.2/10 =.02 molar
now let the degree of dissociation be d
concentration of CO2 =.02d
but Kp=[CO2]
1.16=.02d
d=58%
concentration remain unreacted =42%
Bharat Makkar
9 years ago
sir but the answer is 34%
 
Sunil Kumar FP
9 years ago

sorry for that,i did one mistake
here,Kp=Kc(RT)^delta n
here delta n=1
therefore,Kc=Kp/RT =1.16/.0821*1073 =.01318
Let the degree of dissociation be d
.01318=.02 d

unreated =.02-.01318 =.0068M
interms of mole=.0068*10 =.068 mole
interma of weight =6.8 gm
therefore percentage unreacted =6.8/20)*100 =34%
britto george
9 years ago
this is a simple question and you can get good things from solving this so just try